OpenStudy (anonymous):

A ball is thrown from a height of 217 feet with an initial downward velocity of 17 ft/s h=217-17t-16t^2 Round your answer to the nearest hundreds. Please HELP! <3

OpenStudy (anonymous):

are you solving for time?

OpenStudy (anonymous):

yes it is how long it takes for the ball to hit the ground

OpenStudy (anonymous):

I think you may have wrote your equation wrong: If I'm not mistaken, the kinematic for horizontal freefall is this: Hf-Hi=Vi-1/2gt^2

OpenStudy (anonymous):

\[H_f-H_i=V_i-\frac{ 1 }{ 2 }g*t^2\]

OpenStudy (anonymous):

Thank you so much for helping me. I think I did write it wrong. This little sign I used "^" was reffering to a squared root. The equation to use is t=-b + (square root) b^(exponent 2)-4ac/2x5

OpenStudy (anonymous):

oops I mean ^ as an exponent of 2

OpenStudy (anonymous):

now, just plug into the quadratic and you're are done.

OpenStudy (anonymous):

ive done it so many times and can not figure out the right answer :/ I know you just plug all of it in and then you should get the answer but it keeps coming up as wrong

OpenStudy (anonymous):

\[0=217-17t-16t^2 \rightarrow x= 3.19s\]

OpenStudy (anonymous):

wow you are amazing

OpenStudy (anonymous):

er it should have been t=3.19s but you know what I mean. All I did was plug into a quadratic solver.

OpenStudy (anonymous):

Just have to be careful with how you write the equation out and you'll be fine. good luck.

OpenStudy (anonymous):

Thankyou what is the quadratic solver?

OpenStudy (anonymous):

just google quadratic solver and you will get sites that allow you to plug in the values of a,b, and c and it will spit out your answers.

OpenStudy (anonymous):

I cant even find out how to get to that point to type it in..

OpenStudy (anonymous):

what point?